OpenStudy (anonymous):

f(x) = { x^2; x<=2 { mx+b; x> 2 Find the values of m and b that make the function differentiable and continuous everywhere

OpenStudy (agreene):

Since mx+b is a linear function, \[ m= \mathbb{R},b= \mathbb{R}\]

OpenStudy (anonymous):

Huh...?

OpenStudy (anonymous):

Am I supposed to set the 2 equations equal to each other? (2)^2=2m+b

OpenStudy (agreene):

m = anything b = anything You cant put any constants in there to make them not continuous because it is a linear equation.

OpenStudy (anonymous):

But there is an answer... not just 'anything'... except I do not know how to get it.

OpenStudy (anonymous):

4=2m+b

OpenStudy (anonymous):

at x=2, dy/dx= 2x = 4

OpenStudy (anonymous):

SO m=4

OpenStudy (anonymous):

4=2(4)+b

OpenStudy (anonymous):

b= -4

OpenStudy (anonymous):

You are finding the derivative of 4=2m+b?

OpenStudy (anonymous):

no, the deriv ative of y=x^2, and equating it to the gradient of the line.

OpenStudy (anonymous):

Why are you setting the derivative of y=x^2 = 4?

OpenStudy (anonymous):

Nvm, I understand it now :D Thanks